Find the value of a:
The line through (-2,a) and (a,3) has slope -1/2

Answer by lwsshak3(11628)

Find the value of a:
The line through (-2,a) and (a,3) has slope -1/2
**
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
slope=∆y/∆x
=(3-a)/(a-(-2))
=(3-a)/(a+2)=-1/2
3-a=-a/2-1
a/2=4
a=8